OpenStudy (anonymous):

What is the y-intercept of the line y = 2x + 6?

OpenStudy (anonymous):

plug x=0 so,y=2(0)+6=6

OpenStudy (ash2326):

@elitepenguin did you understand?

OpenStudy (anonymous):

yes thank u very much

OpenStudy (anonymous):

i dont get it really

OpenStudy (anonymous):

6 is the answer coz in such an equation from y=mx+c the m is the gradient while the constant c is the y-intercept so long as you have simplified the equation

OpenStudy (anonymous):

oh i get it now